### v1

@@ 2 Related Works
Sequence-specific DNA-binding proteins are crucial regulators of gene expression in all organisms. Transcription factors (TF) represent a primary example of this class, binding to specific DNA sequences - typically within promoter or enhancer regions - to regulate the transcription of target genes (Spandidos Publications, 2024). By recruiting or blocking RNA polymerase and interacting with various cofactors, TFs precisely coordinate when and where genes are activated or silenced, thus directing essential cellular functions and establishing cell identity. Numerous vital biological processes, such as development, cell cycle control, and responses to environmental stress, rely on sophisticated networks of DNA-binding proteins that act collaboratively. Disruptions or mutations in these proteins can lead to abnormal gene expression, contributing to disease development, including cancers driven by misregulated TFs. In addition, other sequence-specific DNA-binding proteins, such as bacterial restriction enzymes (which recognize and cleave foreign DNA sequences) and DNA repair or recombination proteins (targeting particular DNA motifs), play crucial roles in maintaining genomic integrity. Collectively, sequence-specific DNA binding proteins interpret genomic regulatory information, ensuring precise gene activation and maintaining normal cellular function (Spandidos Publications, 2024). Given their fundamental biological significance, computational methods for identifying and characterizing DNA-binding proteins have attracted extensive research interest. Historically, traditional machine learning approaches employed features such as amino acid composition, conserved motifs, and evolutionary profiles, establishing foundational insights. However, deep learning techniques, particularly convolutional neural networks (CNNs) and recurrent neural networks (RNNs), which take advantage of extensive and comprehensive datasets, have significantly improved the accuracy of DNA-binding residue and motif predictions (IEEE, 2024; Kumar et al., 2024). Integrating protein language models (pLMs), such as ProtTrans, with multi-window CNN architectures has further improved performance, capturing intrinsic biochemical properties and sequence motifs essential for DNA-binding recognition, as demonstrated by a remarkable area under the ROC curve (AUC) of 0.89 (Liu et al., 2024). Furthermore, graph-based neural networks that integrate three-dimensional structural contexts via residue contact maps and spatial graphs have further elevated predictive capabilities. These methods, supported by breakthroughs in protein structure prediction such as AlphaFold2, now offer highly accurate computational tools for genome annotation, elucidation of gene regulatory networks, and the development of targeted gene editing technologies. Advances in protein foundational models not only bolster predictive accuracy, but also open transformative possibilities for the generation of novel DNA-binding proteins. Recent computational design methodologies, exemplified by the work of Glasscock et al., (Glasscock et al., 2024a) utilize powerful generative capabilities inherent in foundational protein models to engineer novel proteins that recognize specific DNA sequences through major groove interactions. These designed proteins exhibited precise target sequence specificity and affinities in the nanomolar range (30–100 nM). Structural validation through crystal structures underscored the high accuracy and reliability of these computational models. Importantly, these newly generated proteins demonstrated effective modulation of transcriptional activity in both Escherichia coli and mammalian cells, underscoring their practical potential. This integration of protein foundational models with computational design significantly advances our ability to create sequence-specific sequence-specific DNA-binding proteins that are easily deployable, highlighting their vast applicability in gene regulation and genome-editing applications (Glasscock et al., 2024a).
@@ 3.3 Protein Binder Generation
Figure 4 provides an example of this process, illustrating the binding interaction between a generated binder and its target. In this case, the target protein $\mathbf{X}$ corresponds to VEGF-A, and the generated binder $\hat{\mathbf{Y}}$ is designed for Chain V. The model learns $p(\mathbf{y}|\mathbf{x})$ and generates $\hat{\mathbf{Y}}$ autoregressively, following the structured input $\mathbf{s}=(\mathbf{x},\text{SEP},\mathbf{y})$. During inference, it conditions on $\mathbf{X}$, initializes decoding with SEP and methionine ($\mathbf{M}$), and samples $\mathbf{y}_{i}$ using: $p_{\text{sample}}(\mathbf{y}_{i}|\mathbf{x},\text{SEP},\mathbf{y}_{<i})$. The structural views in Figure 4 show $\hat{\mathbf{Y}}$ binding to $\mathbf{X}$, with binding affinity $A(\hat{\mathbf{Y}},\mathbf{X})$ quantifying interaction strength. More results on the generation procedure will be discussed in Section 4.2.
Figure 4: Illustration of the protein binding sequence generation process. Top: Target protein Vascular endothelial growth factor VEGF-A (PDB ID 1bj1) in complex with neutralizing antibody. Bottom: Multiple views and representations of a binding protein (in blue) generated by our Prot42 model for Chain V of VEGF-A. The model conditions on the target sequence $\mathbf{X}$ (VEGF-A), followed by SEP token and initial methionine residue, generating binding sequence $\hat{\mathbf{Y}}$ autoregressively using $p(\mathbf{y}|\mathbf{x})$. The generated binder demonstrates a Binding Affinity of 6.6, showcasing the model’s capability to design proteins with specific binding properties.
@@ 4.2 Protein Binders generation
  • 4hsa [^9] https://www.rcsb.org/3d-view/4HSA (interleukin 17a): Interleukin 17A (IL-17A) is a pro-inflammatory cytokine produced primarily by T helper 17 (Th17) cells. It plays a crucial role in host defense against bacterial and fungal infections by inducing the expression of antimicrobial peptides and inflammatory mediators. IL-17A signals through the IL-17 receptor complex, activating downstream pathways such as NF-$\kappa$B and MAPK, leading to the recruitment of neutrophils and enhanced immune responses. Dysregulation of IL-17A is associated with autoimmune diseases, including psoriasis, rheumatoid arthritis, and inflammatory bowel disease.
@@
For fine-tuning Prot42 to generate protein binders, we utilized the STRING database Szklarczyk et al. (2021), a comprehensive resource for protein-protein interactions. STRING integrates experimental data, computational predictions, and text mining to provide confidence scores for protein interactions across multiple organisms.We applied stringent filtering criteria to extract high-quality binding pairs:
@@
    Only interaction pairs with confidence scores $\geq 90\%$ were selected, ensuring high reliability of the binding relationships.
@@
Maximum sequence length 512 tokens
@@
We utilized the same target proteins and binding hotspots that were validated in AlphaProteo studies. These targets represent diverse therapeutic domains: BHRF1 (an Epstein-Barr virus protein that promotes cancer by inhibiting apoptosis), Sc2RBD (SARS-CoV-2 receptor binding domain, critical for viral entry into host cells), IL-7RA (Interleukin-7 receptor alpha, implicated in leukemia and HIV pathogenesis), PD-L1 (Programmed death-ligand 1, a key target in cancer immunotherapy), TrkA (Tropomyosin receptor kinase A, involved in chronic pain and autoimmune conditions), VEGF-A (Vascular endothelial growth factor, critical in cancer and eye disease progression), and TNFalpha (Tumor necrosis factor alpha, a central mediator in autoimmune disease).
@@
We established a comprehensive evaluation framework to assess the quality of generated protein binders, focusing primarily on binding affinity and interface characteristics. Our analysis utilizes Prodigy (PROtein binDIng enerGY prediction), a well-established tool for quantifying protein-protein interactions based on structural features.
@@
    Binding Free Energy ($\Delta G$): The core metric for assessing binding strength, estimated in kcal/mol using Prodigy’s empirical scoring function. Lower (more negative) values indicate stronger predicted binding, with values below -9.0 kcal/mol typically indicating high-affinity interactions.
@@
    Dissociation Constant ($K_{d}$): Calculated from $\Delta G$ using the relationship:
    $K_{d}=e^{\Delta G/RT}$ (1)
    where $R$ is the gas constant and $T$ is temperature (310K). This metric provides a concentration-dependent measure of binding affinity in molar units, facilitating comparison with experimental binding assays.
@@
Figure 6: Examples of generated protein binders (green molecular surfaces - structures in different colors) to target protein sequences provided in the PDB IDs: 1yo5, 1jj6, 1a1f, 8tac, 1oct and 2o4a.
@@
Figure 7 presents the dissociation constants (Kd, in nM) for the highest-affinity binders generated by our model compared to both in silico and laboratory-validated AlphaProteo binders. Lower values indicate stronger binding.
Figure 7: Comparison of binding affinities (Kd, nM) for top binders against therapeutic targets. Green bars represent Prot42 predictions (in silico), blue bars represent AlphaProteo in silico predictions, and light blue bars show laboratory-validated AlphaProteo binders. Lower values indicate stronger binding. The y-axis uses a logarithmic scale to accommodate the wide range of affinity values.
Our results demonstrate that Prot42 consistently generated binders with comparable or superior predicted binding affinities across all tested targets. Particularly notable results include the following: for IL-7RA, our model generated binders with a predicted Kd of 5.3 nM, comparable to AlphaProteo’s 6.3 nM; for PD-L1, Prot42 predicted binders with Kd of 5.2 nM, outperforming AlphaProteo’s 8.2 nM; for TrkA, our model achieved a remarkable Kd of 4.4 nM, significantly better than AlphaProteo’s 13.2 nM; and for VEGF-A, Prot42 generated binders with Kd of 6.6 nM, substantially stronger than AlphaProteo’s 13.4 nM.
Notably, for targets where AlphaProteo has laboratory-validated results, our in silico predictions suggest comparable or better binding. For example, with IL-7RA, our predicted Kd of 5.3 nM compares favorably with AlphaProteo’s laboratory-validated Kd of 0.1 nM, especially considering the typical discrepancy between computational predictions and experimental measurements.
Figure 8: Zooming on the generated protein binder for the RBD of SARS-CoV-2 (PDB ID 6m0j) with the contact region between both proteins and Target hotspot residues.
Figure 8 illustrates our top-ranked generated binder for the receptor-binding domain (RBD) of the SARS-CoV-2 spike glycoprotein (PDB ID 6m0j). For this target, we utilized residues 333-526 of chain E, which encompasses the complete RBD domain responsible for host cell recognition. We conditioned our Prot42 model to generate a binding protein with a minimum length of 50 amino acids to ensure sufficient interface area for stable interaction. The generated binder was specifically filtered from a candidate pool based on proximity (6-8 Å) to critical hotspot residues E485, E489, E494, E500 and E505 which constitute the receptor-binding motif (RBM). These residues are known to form critical interactions with the peptidase domain of human angiotensin-converting enzyme 2 (ACE2). By targeting this specific epitope, our designed binder is predicted to function as a competitive inhibitor, effectively blocking the interaction between the SARS-CoV-2 spike protein and ACE2, thus potentially neutralizing viral entry into host cells.

### v2

@@ 2 Related Works
Sequence-specific DNA-binding proteins are crucial regulators of gene expression in all organisms. Transcription factors (TF) represent a primary example of this class, binding to specific DNA sequences - typically within promoter or enhancer regions - to regulate the transcription of target genes (Spandidos Publications, 2024). By recruiting or blocking RNA polymerase and interacting with various cofactors, TFs precisely coordinate when and where genes are activated or silenced, thus directing essential cellular functions and establishing cell identity. Numerous vital biological processes, such as development, cell cycle control, and responses to environmental stress, rely on sophisticated networks of DNA-binding proteins that often act collaboratively. Disruptions or mutations in these proteins can lead to abnormal gene expression, contributing to disease development, including cancers driven by misregulated TFs. In addition, other sequence-specific DNA-binding proteins, such as bacterial restriction enzymes (which recognize and cleave foreign DNA sequences) and DNA repair or recombination proteins (targeting particular DNA motifs), play crucial roles in maintaining genomic integrity. Collectively, sequence-specific DNA binding proteins interpret genomic regulatory information, ensuring precise gene activation and maintaining normal cellular function (Spandidos Publications, 2024). Given their fundamental biological significance, computational methods for identifying and characterizing DNA-binding proteins have attracted extensive research interest. Historically, traditional machine learning approaches employed features such as amino acid composition, conserved motifs, and evolutionary profiles, establishing foundational insights. However, deep learning techniques, particularly convolutional neural networks (CNNs) and recurrent neural networks (RNNs), which take advantage of extensive and comprehensive datasets, have significantly improved the accuracy of DNA-binding residue and motif predictions (IEEE, 2024; Kumar et al., 2024). Integrating protein language models (pLMs), such as ProtTrans, with multi-window CNN architectures has further improved performance, capturing intrinsic biochemical properties and sequence motifs essential for DNA-binding recognition, as demonstrated by a remarkable area under the ROC curve (AUC) of 0.89 (Liu et al., 2024). Furthermore, graph-based neural networks that integrate three-dimensional structural contexts via residue contact maps and spatial graphs have further elevated predictive capabilities. These methods, supported by breakthroughs in protein structure prediction such as AlphaFold2, now offer highly accurate computational tools for genome annotation, elucidation of gene regulatory networks, and the development of targeted gene editing technologies. Advances in protein foundational models not only bolster predictive accuracy, but also open transformative possibilities for the generation of novel DNA-binding proteins. Recent computational design methodologies, exemplified by the work of Glasscock et al., (Glasscock et al., 2024a) utilize powerful generative capabilities inherent in foundational protein models to engineer novel proteins that recognize specific DNA sequences through major groove interactions. These designed proteins exhibited precise target sequence specificity and affinities in the nanomolar range (30–100 nM). Structural validation through crystal structures underscored the high accuracy and reliability of these computational models. Importantly, these newly generated proteins demonstrated effective modulation of transcriptional activity in both Escherichia coli and mammalian cells, underscoring their practical potential. This integration of protein foundational models with computational design significantly advances our ability to create sequence-specific DNA-binding proteins that are easily deployable, highlighting their vast applicability in gene regulation and genome-editing applications (Glasscock et al., 2024a).
@@ 3.3 Protein Binder Generation
Figure 4 provides an example of this process, illustrating the binding interaction between a generated binder and its target. In this case, the target protein $\mathbf{X}$ corresponds to VEGF-A, and the generated binder $\hat{\mathbf{Y}}$ is designed for Chain V. The model learns $p(\mathbf{y}|\mathbf{x})$ and generates $\hat{\mathbf{Y}}$ autoregressively, following the structured input $\mathbf{s}=(\mathbf{x},\text{SEP},\mathbf{y})$. During inference, it conditions on $\mathbf{X}$, initializes decoding with SEP and methionine ($\mathbf{M}$), and samples $\mathbf{y}_{i}$ using: $p_{\text{sample}}(\mathbf{y}_{i}|\mathbf{x},\text{SEP},\mathbf{y}_{<i})$. The structural views in Figure 4 show $\hat{\mathbf{Y}}$ binding to $\mathbf{X}$, with dissociation constant $K_{d}(\hat{\mathbf{Y}},\mathbf{X})$ quantifying interaction strength. More results on the generation procedure will be discussed in Section 4.2.
Figure 4: Illustration of the protein binding sequence generation process. Top: Target protein Vascular endothelial growth factor VEGF-A (PDB ID 1bj1) in complex with neutralizing antibody. Bottom: Multiple views and representations of a binding protein (in blue) generated by our Prot42 model for Chain V of VEGF-A. The model conditions on the target sequence $\mathbf{X}$ (VEGF-A), followed by SEP token and initial methionine residue, generating binding sequence $\hat{\mathbf{Y}}$ autoregressively using $p(\mathbf{y}|\mathbf{x})$. The generated binder demonstrates a $K_{d}$ of 4.2nM, showcasing the model’s capability to design proteins with specific binding properties.
@@ 4.2 Protein Binders generation
  • 1bj1999https://www.rcsb.org/structure/1BJ1 (vascular endothelial growth factor A): Vascular endothelial growth factor A (VEGF-A) is a potent angiogenic cytokine involved in the regulation of vascular growth and permeability. It exerts its biological functions by binding to VEGF receptors (primarily VEGFR-2) on endothelial cells, initiating downstream signaling cascades that promote endothelial cell proliferation, migration, and survival. The crystal structure (PDB ID: 1BJ1) reveals a homodimeric protein with a characteristic cystine-knot motif, essential for receptor binding. Dysregulated VEGF-A expression contributes to pathological angiogenesis in cancer, diabetic retinopathy, and other diseases, making it a major target for anti-angiogenic therapies.
@@
For fine-tuning Prot42 to generate protein binders, we used the STRING database Szklarczyk et al. (2021), a comprehensive resource for protein-protein interactions. STRING integrates experimental data, computational predictions, and text mining to provide confidence scores for protein interactions across multiple organisms. We applied stringent filtering criteria to extract high-quality binding pairs.
@@
    Only interaction pairs with confidence scores $\geq 90\%$ were selected, ensuring a high reliability of the binding relationships.
@@
Maximum sequence length 512 tokens (including both target and binder)
@@
[…]
@@
We established a comprehensive evaluation framework to assess the quality of generated protein binders, focusing primarily on binding affinity by estimating the Dissociation Constant ($K_{d}$) and interface characteristics. Our analysis utilizes Prodigy (PROtein binDIng enerGY prediction) Vangone A (2015), a well-established tool for quantifying protein-protein interactions based on structural features.
@@
    Dissociation Constant ($K_{d}$): A concentration‐based measure of binding affinity in molar units, defined at equilibrium by
    $K_{d}\;=\;\frac{[P]\,[L]}{[PL]}$
    where $[P]$, $[L]$, and $[PL]$ are the equilibrium concentrations of free protein, free ligand, and protein–ligand complex, respectively. Lower $K_{d}$ values correspond to tighter binding (e.g. $K_{d}<10^{-9}\,\mathrm{M}$ indicates high affinity).
@@
    Binding Free Energy ($\Delta G$): The thermodynamic potential of binding, estimated in kcal/mol (e.g. via Prodigy’s empirical scoring function). It is related to $K_{d}$ by
    $\Delta G\;=\;RT\,\ln K_{d}$ (1)
    where $R$ is the gas constant and $T$ the absolute temperature (typically 310 K). More negative $\Delta G$ denotes stronger binding, with values below about –9.0 kcal/mol generally indicating high‐affinity interactions.
@@
Figure 6: Examples of generated protein binders (green molecular surfaces - structures in different colors) to target protein sequences provided in the PDB IDs.
@@
[…]
Figure 7: Comparison of dissociation constants ($K_{d}$, nM) for top binders against seven therapeutic targets. Green bars represent Prot42 Prodigy predictions; dark‐blue bars, AlphaProteo in silico Prodigy predictions; light‐blue bars, AlphaProteo laboratory measurements; and gray bars, other design methods. Lower $K_{d}$ denotes tighter binding. All values are plotted on a base-10 logarithmic axis.
[…]
Figure 8: Zooming on the generated protein binder for the RBD of SARS-CoV-2 (PDB ID 6m0j) with the contact region between both proteins.
Figure 8 illustrates our generated top-ranked binder for the receptor-binding domain (RBD) of the SARS-CoV-2 spike glycoprotein (PDB ID 6m0j). For this target, we used residues 333-526 of chain E, which encompass the entire RBD domain responsible for host cell recognition. We conditioned our Prot42 model to generate a binding protein with a minimum length of 50 amino acids to ensure sufficient interface area for stable interaction. The generated binder was specifically filtered from a candidate pool based on proximity (6-8 Å) to the critical hotspot residues E485, E489, E494, E500 and E505 which constitute the receptor-binding motif (RBM). These residues are known to form critical interactions with the peptidase domain of human angiotensin-converting enzyme 2 (ACE2). By targeting this specific epitope, our designed binder is predicted to function as a competitive inhibitor, effectively blocking the interaction between the SARS-CoV-2 spike protein and ACE2, thus potentially neutralizing viral entry into host cells.